OpenStudy (anonymous):

Easy Question (Algebra 1) MEDAL FOR WHOEVER HELPS!

OpenStudy (anonymous):

\[\frac{ 3x }{5 } = \frac{ 4 }{ 3 }\]

OpenStudy (zpupster):

cross multiply 3(3x) = 5(4) 9x = 20 divide both sides by 9

OpenStudy (anonymous):

(first day of school)Im doing Honors geography and algebra 2, and like most others im rusty. Forgot how to solve. We are having a review over algebra 1 :p

OpenStudy (anonymous):

x = 20/9 correct?